OpenStudy (anonymous):

Phillip flips an unfair coin eight times. The coin is twice as likely to come up heads as tails. How many times as likely is Phillip to get exactly three heads than exactly two heads?

OpenStudy (anonymous):

step one if the coin is twice as likely to come up heads, what is the probability it comes up heads?

OpenStudy (anonymous):

2/3 to come up with heads.

OpenStudy (anonymous):

ok good now we can use binomial probablility \[P(x=k)=\binom{n}{k}p^k(1-p)^{n-k}\] with \[n=8, k=2, p=\frac{2}{3}, 1-p=\frac{1}{3}\]

OpenStudy (anonymous):

if this is greek i can explain it more slowly but it may look familiar

OpenStudy (anonymous):

I'm still following

OpenStudy (anonymous):

ok so lets compute the probability you get exactly 3 heads it is \[P(x=3)=\binom{8}{3}(\frac{2}{3})^3(\frac{1}{3})^5\]

OpenStudy (anonymous):

a little calculator help gives me \[\frac{448}{6561}\] http://www.wolframalpha.com/input/?i=%288+choose+3%29%282%2F3%29^3%281%2F3%29^5

OpenStudy (anonymous):

oh no that is not the question

OpenStudy (anonymous):

the question is "How many times as likely is Phillip to get exactly three heads than exactly two heads?"

OpenStudy (anonymous):

so we still have to compute \[P(x=2)\]

OpenStudy (anonymous):

let me know when you get that one

OpenStudy (anonymous):

\[\frac{ 112 }{ 6561 }\]

OpenStudy (anonymous):

then divide the first one by the second one to get your answer

OpenStudy (anonymous):

4

OpenStudy (anonymous):

pretty clear that the answer is 4, and it would probably have been easier if we didn’t compute anything, just wrote the two formulas down and cancelled our brains out

OpenStudy (anonymous):

but that is by hindsight and i have no desire to write it down and see you might though

OpenStudy (anonymous):

Something like this? http://gyazo.com/529796e0aeadc0ec583e4fd7fbaaaa90

OpenStudy (anonymous):

\[\frac{\dbinom{8}{3}\left(\frac{2}{3}\right)^3\left(\frac{1}{3}\right)^5}{\dbinom{8}{2}\left(\frac{2}{3}\right)^2\left(\frac{1}{3}\right)^6}\]

OpenStudy (anonymous):

look at all the common factors!

OpenStudy (anonymous):

\[\frac{\binom{8}{3}}{\binom{8}{2}}=2\]

OpenStudy (anonymous):

and also \[\frac{\frac{2}{3}}{\frac{1}{3}}=2\]

OpenStudy (anonymous):

impress your teacher if you have to hand it in do it the smart way